OpenStudy (smartballerina01):

A quadratic equation is shown below: x2 + 5x + 4 = 0 Part A: Describe the solution(s) to the equation by just determining radicand. Show your work. (3 points) Part B: Solve 4x2 -12x + 5 = 0 using an appropriate method. Show the steps of your work, and explain why you chose the method used. (4 points) Part C: Solve 2x2 -10x + 3 = 0 by using a method different from the one you used in Part B. Show the steps of your work. (3 points)

jhonyy9 (jhonyy9):

x^2 +5x +4 =0 (x+4)(x+1) =0

OpenStudy (wolf1728):

a=1 b=5 c=4 the radicand is square root of b^2 - 4*a*c

OpenStudy (wolf1728):

smartballerina01 do you want to finish solving the radicand?

OpenStudy (zzr0ck3r):

Weird. The radicand is what is under a radical. I think they mean the discriminant?

OpenStudy (welshfella):

Ive never heard of a radicand..

OpenStudy (smartballerina01):

I think they mean discriminant

OpenStudy (welshfella):

Part B: this can be factored using the ac method would work here

OpenStudy (welshfella):

ac is first coefficient times last = 20. -2 * -10 = 20 so try these : 4x^2 - 12 x+ 5 = 0 4x^2 - 10x - 2x + 5 = 0 4x^2 - 2x - 10x + 5 = 0 factor by grouping 2x(2x - 1) - 5(2x - 1) = 0 now take out 2x - 1 can you continue?

OpenStudy (welshfella):

if you take out 2x - 1 what is left?